x/6 - x/8 = 1 In this problem I have to solve and check for x. I tried the following:
x/6 4/24
-
x/8 3/24

1/24
Is this correct or did I do something wrong?

Answer by checkley71(8403)

X/6-X/8=1
(4X-3X)/24=1
X/24=1
X=1*24
X=24
-------------------------
YOU LOST THE X FACTOR.
------------------------
PROOF
24/6-24/8=1
4-3=1
1=1